OpenStudy (anonymous):

Given the replacement set {2, 3, 4, 5}, solve 6x – 2 = 5x + 3. A. 4 B. 3 C. 5 D. 2

OpenStudy (anonymous):

@izuru

OpenStudy (anonymous):

Plug all the values as x and see if left hand = right hand

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

Nope, try again

OpenStudy (anonymous):

dang

OpenStudy (anonymous):

So this is what I mean, you tried 4, 6(4)-2 = 5(4)+3 22=23 that isn't right :P

OpenStudy (anonymous):

So try the other ones in the set given

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

no its C

OpenStudy (anonymous):

Yeah there you go :)

OpenStudy (anonymous):

\[6(5)-2=5(5) + 3 \implies 30-2=25+3 \implies 28=28 \checkmark\]